Suppose you want to have $300,000 for retirement in 25 years. Your account earns 10% interest. Round your answers to the nearest cent.
a) How much would you need to deposit in the account each month?
$
b) How much interest will you earn?
$

Answer by Boreal(15235)

300000=m(1+(r/n))^300-1)/(r/n), where r=0.10 and n=12
multiply through by (0.10/12)
2500=m(1.0083333)^300-1=m(11.05694,) but round only at the end.
m=$226.10
-
The amount of all the deposits is $226.10*300=$67,830
The interest is everything else up to $300,000=$232,170